# Visual Imitation Enables Contextual Humanoid Control

### v4

@@ 3.1 Preprocessing
We preprocess monocular RGB videos with off-the-shelf state-of-the-art human pose estimation and SfM methods. First, people are detected and associated across frames using Grounded SAM2 [45, 46]. For each detected person, we recover per‑frame 3D SMPL [24] parameters with VIMO [47], obtaining per‑frame local pose $\theta^{t}$, shape $\beta$, and SMPL 3D joints $J_{\mathrm{3D}}^{\,t}\in\mathbb{R}^{J\times 3}$. We detect 2D keypoints $J_{\mathrm{2D}}^{\,t}$, i.e., body joint pixel positions, with ViTPose [48]. Foot contact is regressed by BSTRO [49]. For scene reconstruction, we obtain the world point cloud from either MegaSaM [35] or MonST3R [36], which is parameterized as per-frame depth $D^{t}$, camera pose $[R^{t}|t^{t}]$, and a shared camera intrinsic matrix $K$. Note that the resulting point cloud is not metrically accurate.
@@ 4 Policy Learning
Observations. Our policies are conditioned on both proprioceptive and target-related observations. The proprioceptive inputs include a history of the robot’s joint positions ($q$), joint velocities ($\dot{q}$), angular velocity ($\omega$), projected gravity vector ($g$), and previous actions ($a^{t-1}$); we use a history length of 5 in practice. In addition, the policy receives local target observations: the target joint angles, target root roll and pitch, and the desired root direction, specified by relative x-y offset and yaw angle between the robot’s current root position and the target root, all expressed in the robot’s local frame. For policies conditioned on heightmaps, we further provide an elevation map around the torso. This is represented as an $11\times 1$ 11 grid sampled at $0.1\mathrm{m}$ intervals, which captures local terrain geometry. Finally, the critic receives additional privileged observations, which are detailed in Table 3.
@@ 5.1 Reconstruction and Data
Versatility. Figure 4 highlights the breadth of our reconstruction pipeline, showcasing (i) robust environment reconstruction from an Internet video involving dynamic human-scene interaction, (ii) multi-human reconstruction and retargeting. Furthermore, the dense point cloud reconstruction enables ego-view RGB-D rendering via simple rasterization. While not used in our current policy, this offers a promising direction for future work—especially given the challenges of rendering naturalistic images in simulation.
@@ 5.2 Real-world Deployment
Setup. We deploy our controller on a 23-DoF Unitree G1 humanoid and run it onboard at 50 Hz. Following [61], we set relatively low joint gains, $K_{p}=75$, to avoid excessively fast or overly stiff behaviour—which helps to avoid excessively violent contact when the robot makes heavy contact with objects such as chairs or stairs. Height-maps are computed in real time using Fast-lio2 [62] and probabilistic terrain mapping [63, 17]. We feed joystick targets from a human operator. Including policy running, all operations are run onboard. We found two critical ingredients for successful motion deployment through iterative sim-to-real trials: (i) relaxing the episode-termination tolerances with respect to the reference motion, and (ii) injecting realistic physics perturbations during training. Full details are given in Sec. C.
@@ 7 Limitations
At test time, the controller receives only proprioception and an $11\times 1$ 11 LiDAR height‑map. This coarse grid is adequate for terrain and chairs but lacks the resolution for precise contacts, manipulation, or reasoning about overhanging obstacles. Incorporating richer perceptual inputs—such as RGB‑D data or learned occupancy grids—would likely broaden the method’s applicability and improve its semantic understanding of the environment.
@@ A.1 Notations
$x_{\text{2D}}=\Pi\left(K\begin{bmatrix}R&t\end{bmatrix}\begin{bmatrix}x_{\text{3D}}\\ 1\end{bmatrix}\right),$ (2)
@@ A.3 Generating Simulation Ready Data
Gravity alignment. We first estimate the gravity direction in the initial camera frame and define the transformation that converts the 3D reconstruction from world coordinates to a gravity-aligned coordinate system compatible with physics engines. GeoCalib [51] provides the roll–pitch angles for that frame; the composite rotation $R_{\text{gravity,world}}\!=\!R_{y\uparrow\!\to z}R_{x}(\pi)R_{z}(\rho)R_{\text{cam,world}}$ re-orients the reconstruction so that $+z$ points up, ensuring consistency with gravity in physics engines. The transformation is applied to both human keypoints and static scene geometry.
Pointcloud filtering. For each world pointcloud we discard noisy background points and human pointclouds with thresholding on depth gradient, rotate by $R_{\text{gravity,world}}$, scale, and crop to a ±$\pm 2\mathrm{\,m}$ box around the SMPL joints per frame. A $0.1\mathrm{\,m}$ voxel grid then keeps at most 20 samples per occupied cell, shrinking the pointcloud to about 5% of its original size without losing surface detail.
@@ B.1 RL Setup and Training
We use the PPO algorithm [53]. We use a discount factor of $\gamma=0.99$, and a GAE parameter of $\lambda=0.95$. We use an adaptive learning rate with a desired KL of $0.02$. For adaptive learning rate, we use $1.2$ as opposed to the usual $1.5$ learning rate change, as we find that this leads to faster training. When doing RL from scratch, we set the learning rate (which is modified from this value by adaptive KL) at the start to $31e-$ 3 however when finetuning we find it is important to start from a very low learning rate of $52e-$ 5 to prevent early updates with a high learning rate from destroying the existing checkpoint when the distribution of data is biased early during an episode when all environments start. We use a max gradient norm of 1.0. 5 learning epochs are used per rollout, and our rollout length is 24. Our entropy coefficient is set to $0.0025$. We use 2 x NVIDIA 4090 GPUs for training, each with 4096 environments (so an effective total of 8192 environments). We use MLPs with 4 layers of [1024, 512, 256, 128] dimensions for different layers.
@@ B.2 Observations
Base linear velocity (last 5) $5\times 3=15$D
@@ B.5 Reward Formulation
Action‐Rate Penalty $\lVert\mathbf{a}_{t}-\mathbf{a}_{t-1}\rVert^{2}$ $-8.0$ Can’t move too quickly on real robot
Ankle Action Penalty $\lVert\mathbf{a}_{t}\rVert^{2}$ $-4.0$ Ankles should stay near neutral
[…]
Collision Penalty $\sum 1\!\bigl{[}\lVert\mathbf{f}_{c}\rVert>0.1\bigr{]}$ $-1.0$ Discourage collisions on penalized links
Contact–No–Velocity Pen. $\lVert\mathbf{v}_{\text{feet}}\rVert_{(c>1)}$ $-100$ Penalize stationary foot contact
@@
Link Vel Tracking $\exp\!\bigl{(}-k\,\lVert\dot{\mathbf{p}}_{\text{links}}-\dot{\mathbf{p}}^{\ast}\rVert^{2}\bigr{)}$ $5$ $0.1$ Track link velocities
Feet Contact Match $\sum\mathbbm{1}\!\bigl{[}c_{t}^{(i)}=c_{t}^{\ast(i)}\bigr{]}$ $1$ Match contacts from reference motion (BSTRO)
@@
Termination Penalty episode end (failure) $-500$ Harsh penalty on failure events
@@ B.7 Domain Randomization
[…]
@@
Gravity bias $\Delta g$ $0.01$g Fixed bias added to gravity vector for an episode in policy input.
DoF-position bias $\Delta q$ $0.005$rad Fixed bias added to each joint position in policy input.

### v5

@@ 3.1 Preprocessing
We preprocess monocular RGB videos with off-the-shelf state-of-the-art human pose estimation and SfM methods. First, people are detected and associated across frames using Grounded SAM2 [45, 46]. For each detected person, we recover per‑frame 3D SMPL [24] parameters with VIMO [47], obtaining per‑frame local pose $\theta^{t}$, shape $\beta$, and SMPL 3D joints $J_{\mathrm{3D}}^{\,t}\in\mathbb{R}^{J\times 3}$. We detect 2D keypoints $J_{\mathrm{2D}}^{\,t}$, i.e., body joint pixel positions, with ViTPose [48]. Foot contact is regressed by BSTRO [49]. For scene reconstruction, we obtain the world point cloud from either MegaSaM [35] or MonST3R [36], which is parameterized as per-frame depth $D^{t}$, camera pose $R^{t}|t^{t}]$, and a shared camera intrinsic matrix $K$. Note that the resulting point cloud is not metrically accurate.
@@ 4 Policy Learning
Observations. Our policies are conditioned on both proprioceptive and target-related observations. The proprioceptive inputs include a history of the robot’s joint positions ($q$), joint velocities ($\dot{q}$), angular velocity ($\omega$), projected gravity vector ($g$), and previous actions ($a^{t-1}$); we use a history length of 5 in practice. In addition, the policy receives local target observations: the target joint angles, target root roll and pitch, and the desired root direction, specified by relative x-y offset and yaw angle between the robot’s current root position and the target root, all expressed in the robot’s local frame. For policies conditioned on heightmaps, we further provide an elevation map around the torso. This is represented as an $11\times 11$ grid sampled at $0.1\mathrm{m}$ intervals, which captures local terrain geometry. Finally, the critic receives additional privileged observations, which are detailed in Table 3.
@@ 5.1 Reconstruction and Data
Versatility. Figure 4 and our webpage’s Viser [61] visualizations highlight the breadth of our reconstruction pipeline, showcasing (i) robust environment reconstruction from an Internet video involving dynamic human-scene interaction, (ii) multi-human reconstruction and retargeting. Furthermore, the dense point cloud reconstruction enables ego-view RGB-D rendering via simple rasterization. While not used in our current policy, this offers a promising direction for future work—especially given the challenges of rendering naturalistic images in simulation.
@@ 5.2 Real-world Deployment
Setup. We deploy our controller on a 23-DoF Unitree G1 humanoid and run it onboard at 50 Hz. Following [62], we set relatively low joint gains, $K_{p}=75$, to avoid excessively fast or overly stiff behaviour—which helps to avoid excessively violent contact when the robot makes heavy contact with objects such as chairs or stairs. LiDAR odometry and height-maps are computed in real time using Fast-lio2 [63] and probabilistic terrain mapping [64, 17]. The odometry is used to align the local height-map observations into the robot’s frame. We feed joystick targets from a human operator. Including policy running, all operations are run onboard. We found two critical ingredients for successful motion deployment through iterative sim-to-real trials: (i) relaxing the episode-termination tolerances with respect to the reference motion, and (ii) injecting realistic physics perturbations during training. Full details are given in Sec. C.
@@ 7 Limitations
At test time, the controller receives only proprioception and an $11\times 11$ LiDAR height‑map. This coarse grid is adequate for terrain and chairs but lacks the resolution for precise contacts, manipulation, or reasoning about overhanging obstacles. Incorporating richer perceptual inputs—such as RGB‑D data or learned occupancy grids—would likely broaden the method’s applicability and improve its semantic understanding of the environment.
@@ A.1 Notations
[…]
@@ A.3 Generating Simulation Ready Data
Gravity alignment. We first estimate the gravity direction in the initial camera frame and define the transformation that converts the 3D reconstruction from world coordinates to a gravity-aligned coordinate system compatible with physics engines. GeoCalib [51] provides the roll–pitch angles for that frame; the composite rotation $R_{\text{gravity,world}}\!=\!R_{y\uparrow\!\to z}R_{x}(\pi)R_{z}(\rho)R_{\text{cam,world}}$ re-orients the reconstruction so that +z+z points up, ensuring consistency with gravity in physics engines. The transformation is applied to both human keypoints and static scene geometry.
Pointcloud filtering. For each world pointcloud we discard noisy background points and human pointclouds with thresholding on depth gradient, rotate by $R_{\text{gravity,world}}$, scale, and crop to a $\pm 2\mathrm{\,m}$ box around the SMPL joints per frame. A $0.1\mathrm{\,m}$ voxel grid then keeps at most 20 samples per occupied cell, shrinking the pointcloud to about 5% of its original size without losing surface detail.
@@ B.1 RL Setup and Training
We use the PPO algorithm [53]. We use a discount factor of $\gamma=0.99$, and a GAE parameter of $\lambda=0.95$. We use an adaptive learning rate with a desired KL of $0.02$. For adaptive learning rate, we use $1.2$ as opposed to the usual $1.5$ learning rate change, as we find that this leads to faster training. When doing RL from scratch, we set the learning rate (which is modified from this value by adaptive KL) at the start to 1e31e-3 however when finetuning we find it is important to start from a very low learning rate of 2e52e-5 to prevent early updates with a high learning rate from destroying the existing checkpoint when the distribution of data is biased early during an episode when all environments start. We use a max gradient norm of 1.0. 5 learning epochs are used per rollout, and our rollout length is 24. Our entropy coefficient is set to $0.0025$. We use 2 x NVIDIA 4090 GPUs for training, each with 4096 environments (so an effective total of 8192 environments). We use MLPs with 4 layers of [1024, 512, 256, 128] dimensions for different layers.
@@ B.2 Observations
Fast-lio2 [63] and probabilistic terrain mapping [64, 17] are used to derive heightmaps and a live odometry signal in real. This can be used to follow 2D commands in the global frame as we did in testing (see Figure 9) and as is done in a concurrent work [69] for the purpose of teleoperation. However, we note that we do not do this in the final version of the paper to avoid any hysteresis in the position commands, and instead multiply the joystick value by a fixed amount to obtain a position which is interpreted by the policy as an offset to the target position in the local frame.
@@
Base linear velocity (last 5) $5\times 3=15$D $\times$ $\times$ $\times$
@@ B.5 Reward Formulation
Action‐Rate Penalty $\lVert\mathbf{a}_{t}-\mathbf{a}_{t-1}\rVert^{2}$ 8.0-8.0 Can’t move too quickly on real robot
Ankle Action Penalty $\lVert\mathbf{a}_{t}\rVert^{2}$ 4.0-4.0 Ankles should stay near neutral
DOF-Pos Limits Penalty $\sum[\text{overflow w.r.t.\;0.98\,range}]$ 50.0-50.0 Penalize joints outside 98 % of their range
Collision Penalty $\sum 1\!\bigl{[}\lVert\mathbf{f}_{c}\rVert>0.1\bigr{]}$ 1.0-1.0 Discourage collisions on penalized links
Contact–No–Velocity Pen. $\lVert\mathbf{v}_{\text{feet}}\rVert_{(c>1)}$ 100-100 Penalize stationary foot contact
@@
Link Vel Tracking $\exp\!\bigl{(}-k\,\lVert\dot{\mathbf{p}}_{\text{links}}-\dot{\mathbf{p}}^{\ast}\rVert^{2}\bigr{)}$ 55 $0.1$ Track link velocities
Feet Contact Match $\sum\mathbbm{1}\!\bigl{[}c_{t}^{(i)}=c_{t}^{\ast(i)}\bigr{]}$ 11 Match contacts from reference motion (BSTRO)
@@
Termination Penalty episode end (failure) 500-500 Harsh penalty on failure events
@@ B.7 Domain Randomization
DoF friction $\mu_{\text{DoF}}$ $\mathcal{U}[0,\,0.02]$ Different per environment.
Random pushes (xy) $\Delta v_{\text{push},xy}$ $\mathcal{U}[-0.25,\,0.25]$ m/s Additive to the robot’s reference state or current velocity.
Random pushes (z) $\Delta v_{\text{push},z}$ $\mathcal{U}[-0.10,\,0.10]$ m/s Additive to the robot’s reference state or current velocity.
@@
Gravity bias $\Delta g$ 0.010.01g Fixed bias added to gravity vector for an episode in policy input.
DoF-position bias $\Delta q$ 0.0050.005rad Fixed bias added to each joint position in policy input.